OpenStudy (anonymous):

If 9x-8y=4 and 2x-3y=-4, then what is the value of x?

OpenStudy (anonymous):

step1; multiply first equation with 3 step2; multiply second equation with -8 step3; add the two equations see what you get

OpenStudy (anonymous):

9x - 8y = 4 ---> (2)9x - 8y = 4 2x - 3y = - 4--> (-9)2x - 3y = - 4 --------------- 18x - 16y = 8 (result of multiplying by 2) -18x + 27y = 36 (result of multiplying by -9) ----------------add 0 + 11y = 44 11y = 44 (divide both sides by 11) y = 4 now sub 4 in for y in either of the original equations 2x - 3y = - 4 2x - 3(4) = - 4 2x - 12 = - 4 (add 12 to both sides) 2x - 12 + 12 = - 4 + 12 (combine like terms) 2x = 8 (divide by 2 to get x by itself) 2x/2 = 8/2 x = 4 now check by subbing in known values into either of the original equations 9x - 8y = 4 9(4) - 8(4) = 4 36 - 32 = 4 4 = 4 (correct) ANSWER : x = 4, y = 4 or (4,4) Do you have any questions ?